# MIRA: Medical Time Series Foundation Model for Real-World Health Data

### v5

@@ 1 Introduction
Medical time series data, including signals such as electrocardiograms (ECG) [1], electroencephalograms (EEG) [2], vital signs and laboratory measurements [3], are key to understanding the dynamic physiological states of patients over time [4, 5]. Continuously modeling these patient data trajectories supports clinical forecasting tasks such as predicting organ failure or treatment response [1, 6], enabling earlier decisions and better patient outcomes [7, 8]. However, effectively utilizing medical time series data in practice remains challenging, as patient populations, disease profiles, and clinical protocols can vary widely across regions and institutions [9]. Moreover, such data is often collected in an irregular and asynchronous manner [10, 11], further complicating efforts to build generalizable models. These challenges are amplified by regulatory frameworks such as GDPR [12, 13], which restrict cross-institutional data exchange and hinder the development of unified modeling paradigms, often resulting in redundant and resource-intensive local optimization efforts. This highlights the need for a foundation model trained on validated medical time series datasets, that is able to learn generalizable patterns from large and varied datasets, reducing the need for extensive data annotation and custom model building, while also helping to share knowledge effectively across different clinical institutions, data types, and medical tasks [14, 15].
@@
However, creating effective medical foundation models for time series is challenging, because clinical data are inherently highly irregular and varied: signals like ECGs recorded at millisecond intervals [16] are found alongside laboratory tests conducted hours apart [17] (Figure 1). Differences in clinical workflows and equipment limits result in irregular sampling, missing values, and inconsistent temporal dependencies [18, 19]. Furthermore, physiological variables differ widely in their value and frequency ranges, requiring models that can handle these varied scales and types of data. Traditional methods using medical task-specific time-series models [20, 21, 22, 23, 24, 25], trained on isolated datasets, often do not generalize well to new situations and are difficult to scale [26, 27, 10]. Meanwhile recent general-purpose time series foundation models [28, 29, 30, 31, 32, 33] show promise, but they usually assume uniform time intervals. Similarly, emerging medical time series foundation models demonstrate cross-dataset generalization but only within narrowly defined domains—for example, across EEG datasets for sleep monitoring or across datasets for Alzheimer’s diagnosis, and cannot handle continuous-time forecasting or irregular sampling [34, 35, 36, 37]. These gaps—where current models struggle with the complex, irregular, and multi-frequency nature of medical data [28, 38]—motivates our research and raises a central question:
@@
To address this central question, we introduce MIRA—a Medical Foundation Model for Irregularity Adaptation. MIRA is specifically designed to address the challenges of medical time series: First, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE), which extends the standard RoPE [39] to handle time values using learnable frequency adjustments, allowing MIRA to accurately model the varying frequency intervals of irregularly sampled data. Second, to further improve its ability to adapt to different time features in medical signals, MIRA uses a frequency-specific mixture-of-experts (MoE) layer that learns to route frequency-related information through specialized components. Finally, MIRA integrates a Continuous Dynamics Extrapolation Block, using Neural Ordinary Differential Equations (Neural ODEs) [40]. This feature allows the model to predict patient health paths at any future time point, moving beyond the limits of fixed observation grids. MIRA is pretrained on a large curated collection of medical time series, comprised of over 454 billion time points from publicly available datasets. These datasets cover readings from intensive care units (ICUs), operating rooms, pediatric critical care, and long-term sleep and mental health monitoring. All data undergo preprocessing that standardizes time alignment, normalizes sampling frequencies, and maintains clinical realism. MIRA outperforms other time series foundation models with a similar number of activated parameters across various real-world benchmarks, achieving reductions in forecasting errors by an average of 8% and 6% in out-of-distribution and in-distribution scenarios, respectively.
@@ 2 Related Work
Medical Time Series Models. Machine learning has enabled significant progress in healthcare time series analysis, supporting a wide range of tasks, including dynamic forecasting [21, 22, 41, 42], survival analysis [43, 44], clustering and phenotyping [45, 46, 47], screening and monitoring [23, 48, 49], early diagnosis [24, 25, 50], pharmacology [51], treatment effect estimation [52, 53, 42], epidemiological and pandemic influenza surveillance [54], and hospital resource allocation [20, 55]. Despite this progress, most existing methods are narrowly designed for specific tasks or datasets, limiting their adaptability to diverse clinical scenarios [56]. Recently, foundation models for medical time series have emerged, showing promising cross-dataset generalization [34, 37, 35]. However, these models focus primarily on classification tasks and are not designed to handle continuous-time forecasting or irregular temporal patterns common in real-world clinical data.
Generalist Time Series Foundation Models. Recent foundation models have shown strong zero-shot forecasting without task-specific tuning [57, 58, 59, 60, 61]. Chronos [33] first adapted T5 [62] by discretizing time series intervals as text tokens. More recent models—Moirai [28], Sundial [32], and TimesFM [30]—improve support for variable dimensions and frequencies. Time-MoE [29] and Moirai-MoE [31] further improve specialization with sparse experts. However, these models largely require regularly sampled data, limiting their use on irregular clinical data. See Appendix A for more information.
Irregular Time Series Models. Modeling irregular time series has attracted increasing attention, particularly for applications requiring continuous-time reasoning. Early works such as Neural ODEs [40, 63] and State Space Models (SSMs)[64, 65] have demonstrated the ability to capture continuous dynamics by parameterizing the evolution of latent states over time. Another line of research focuses on adapting deep neural architectures, such as RNNs [66, 67, 68] and Transformers [19, 69, 70, 71], to irregular settings. While these models have shown success in tasks like interpolation and classification, their effectiveness for long-term forecasting remains underexplored.
@@ 3.1 Irregular Medical Time Series Forecasting
Let each time series instance be represented as paired sequences of timestamps and observations: $\{(t_{i},x_{i})\}_{i=1}^{N}$ where $t_{i}\in\mathbb{R}^{+}$ are strictly increasing ($t_{1}<\cdots<t_{N}$) and $x_{i}\in\mathbb{R}\cup\{\text{NaN}\}$ contains real values or missing entries. Our model is designed to handle two prevalent irregular patterns: (1) Regular-grid missing values. Observations occur at equidistant timestamps $t_{i}=i\Delta t$ but with time-level missing values: some timestamps may have missing measurements. Formally, there exists a subset $\mathcal{M}\subseteq{1,\ldots,N}$ such that the corresponding observations $x_{i}$ are missing (i.e., $x_{i}=\text{NaN}$) for all $i\in\mathcal{M}$, while the timestamps $t_{i}$} remain fully observed. (2) Irregular sampling. Timestamps $t_{i}$ follow non-uniform intervals with $\Delta t_{i}=t_{i+1}-t_{i}$ varying for different $i$. All observations at existing timestamps are present ($x_{i}\in\mathbb{R}$), but the temporal spacing is irregular. Adhering to the channel-independent setting [60], we formulate the problem in a uni-variate time series manner [28]. Given history $\{(x_{i},t_{i})\}_{i=1}^{L}$, the forecasting task requires predicting future values at known target timestamps $\{t_{L+1},...,t_{L+H}\}$:
@@ 3.2.1 Continuous-Time Rotary Positional Encoding
Standard RoPE [39] assumes discrete and uniformly spaced token indices, which limits its applicability to clinical time series with real-valued and irregular timestamps. To address this, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE) that generalizes RoPE to operate directly on continuous-time inputs. Let $t\geq 0$ denote the timestamp of a token in an irregularly sampled sequence, and let $d$ be the model dimensionality, assumed to be even. Specifically, we discretize continuous timestamps $t\geq 0$ into rotation angles without assuming fixed intervals, enabling the model to handle irregular sampling (more detail can be found at Appendix B). The angular frequencies $\{\omega_{i}\}_{i=0}^{d/2-1}$ are defined as: $\omega_{i}=10000^{-2i/d}$. The resulting time-dependent rotation angle is given by:
@@
This implies that the positional interaction between two tokens is solely a function of their relative timestamp offset (nm)(n-m). In CT-RoPE, this property generalizes to real-valued timestamps, enabling the attention mechanism to capture continuous-time relational structure while retaining the efficiency and structure of standard dot-product attention. In line with Chowdhery et al. [72], we remove biases except in QKV projections to enhance extrapolation. More detail can be found in Appendix C.
@@ 3.2.3 Continuous Dynamics Extrapolation Block
Auto-regressive transformer architectures, which generate predictions in a stepwise manner under causal masking, cannot incorporate the timestamp of the target token during inference, as it is not available until after generation. To address this limitation and enable extrapolation to arbitrary timestamps, we introduce a Neural ODE-based [40] extrapolation module that evolves the latent state from the current token’s timestamp to the target prediction token timestamp, allowing time-aware forecasting at unseen or irregular time points.
@@ 3.3 Model Training
To support generalizable and clinically relevant time series modeling, we curate a large pretraining corpus spanning over 454 billion time points from various real-world healthcare settings. The dataset collection includes signals from ICUs a nd operating rooms, pediatric critical care, long-term sleep and mental health monitoring, and population-level epidemiological surveillance. All data are drawn from publicly available clinical datasets, including MIMIC-III [73], MIMIC-IV [74], PTB-XL [75], Sleep-EDF [76], and the WAVES Pediatric Waveform Database [77]. To enable the model to acquire general knowledge, we resumed training from the Time-MoE checkpoint. A full summary of included datasets and the pre-processing applied is provided in Appendix E.
@@
We adopt a similar model configuration strategy following Time-MoE [29] by providing three model variants of increasing scale: $MIRA_{small}$, with approximately 73 million parameters; $MIRA_{base}$, with approximately 114 million parameters; $MIRA_{large}$, with approximately 455 million parameters;All models are trained on max to eight NVIDIA 80G H/A100 GPUs, using a micro batch size of 128 , and a maximum sequence length of 512. For model configurations and training details, refer to Appendix F.
@@ 4.1 Experiment Setup
Downstream Clinical Datasets. We evaluate MIRA on a diverse suite of real-world clinical and public health datasets, spanning multi-modal physiological signal analysis, critical care monitoring, ambulatory bio-signal tracking, epidemiological surveillance, and healthcare resource utilization. To systematically study temporal robustness, we group datasets into two categories: (1) inherently irregular datasets, which exhibit irregular sampling and natural missing values due to clinical or observational workflows. These are CinC 2012 [78] and CinC 2019 [79]. Furthermore, we use (2) originally regular datasets, i.e.,MIT-BIH [80], Johns Hopkins COVID-19 Dataset [81] ,CDC Influenza Hospitalizations Admissions (CDC-IHA)  [^1] https://www.cdc.gov/flu-forecasting/data-vis/current-week.html, Heart Rate [82] and illness [83], for which we simulate irregularity by randomly masking 30% of time points. The full list of datasets and statistics is provided in Appendix G.
Baselines. We compare MIRA against 13 state-of-the-art forecasting models, which we categorize into three groups: (i) zero-shot foundation models, including Time-MoE [29]  [^2] We excluded Time-MoE Ultra and Sundial from baselines as they are not open-sourced., Moirai [28], Moirai-MoE [31], Moment [84], TimeGPT [85], Timer [86], Lag-Llama [87], TimesFM [30], and Chronos [33], which require inputs to be interpolated for evaluation; and (ii) full-shot forecasting models, including ContiFormer [19], T-PatchGNN [88], Neural-CDE [89], and ODE-RNN [66], which are specialized for irregularly sampled time series and require task-specific training. (iii) Continue pre-trained zero-shot foundation models on medical corpora, including Time-MoE [29], Moirai [28] and Chronos [33] which performance best on zero-shot evaluation. Implementation details for all baselines are provided in Appendix H.

### v6

@@ 1 Introduction
Medical time series data, including signals such as electrocardiograms (ECG) [gupta2024comprehensive], electroencephalograms (EEG) [ekambaram2023tsmixer], vital signs and laboratory measurements [van2023cross], are key to understanding the dynamic physiological states of patients over time [DBLP:journals/corr/abs-2504-17613, DBLP:journals/corr/HarutyunyanKKG17]. Continuously modeling these patient data trajectories supports clinical forecasting tasks such as predicting organ failure or treatment response [gupta2024comprehensive, guo2023ehr], enabling earlier decisions and better patient outcomes [DBLP:conf/bcb/LiuZJS22, li2024cif]. However, effectively utilizing medical time series data in practice remains challenging, as patient populations, disease profiles, and clinical protocols can vary widely across regions and institutions [wu2025term2note]. Moreover, such data is often collected in an irregular and asynchronous manner [sahiner2023data, guo2022evaluation], further complicating efforts to build generalizable models. These challenges are amplified by regulatory frameworks such as GDPR [DBLP:conf/clef/SchlegelLW0NKBZ23, li2019impact], which restrict cross-institutional data exchange and hinder the development of unified modeling paradigms, often resulting in redundant and resource-intensive local optimization efforts. This highlights the need for a foundation model trained on validated medical time series datasets, that is able to learn generalizable patterns from large and varied datasets, reducing the need for extensive data annotation and custom model building, while also helping to share knowledge effectively across different clinical institutions, data types, and medical tasks [DBLP:conf/bionlp/LiWSBNKZB0N23, DBLP:journals/corr/abs-2404-03264].
@@
However, creating effective medical foundation models for time series is challenging, because clinical data are inherently highly irregular and varied: signals like ECGs recorded at millisecond intervals [DBLP:conf/iclr/NieNSK23] are found alongside laboratory tests conducted hours apart [pohl2025generating] (Figure 1). Differences in clinical workflows and equipment limits result in irregular sampling, missing values, and inconsistent temporal dependencies [DBLP:conf/mlhc/LiptonKW16, DBLP:conf/nips/ChenRWFSL23]. Furthermore, physiological variables differ widely in their value and frequency ranges, requiring models that can handle these varied scales and types of data. Traditional methods using medical task-specific time-series models [DBLP:journals/titb/WangYPEWHGW22, DBLP:conf/icml/FerteDHGJTLH24, DBLP:conf/ijcai/ChenYW024, DBLP:conf/ijcai/JiaYZDJ24, DBLP:conf/aaai/ZhangZL24, DBLP:conf/ijcai/SohnPPC24], trained on isolated datasets, often do not generalize well to new situations and are difficult to scale [DBLP:journals/access/YangQZ23, DBLP:conf/kdd/LiangWNJ0SPW24, sahiner2023data]. Meanwhile recent general-purpose time series foundation models [DBLP:conf/icml/WooLKXSS24, DBLP:journals/corr/abs-2409-16040, DBLP:conf/icml/DasKSZ24, DBLP:journals/corr/abs-2410-10469, DBLP:journals/corr/abs-2502-00816, DBLP:journals/corr/abs-2403-07815] show promise, but they usually assume uniform time intervals. Similarly, emerging medical time series foundation models demonstrate cross-dataset generalization but only within narrowly defined domains—for example, across EEG datasets for sleep monitoring or across datasets for Alzheimer’s diagnosis, and cannot handle continuous-time forecasting or irregular sampling [DBLP:journals/corr/abs-2410-03794, DBLP:journals/npjdm/WornowXTPSFPFS23, DBLP:journals/npjdm/GuoFSFMAPSS24, DBLP:conf/icml/ThapaHKMGM024]. These gaps—where current models struggle with the complex, irregular, and multi-frequency nature of medical data [DBLP:conf/icml/WooLKXSS24, DBLP:conf/iclr/OreshkinCCB20]—motivates our research and raises a central question:
@@
To address this central question, we introduce MIRA—a Medical Foundation Model for Irregularity Adaptation. MIRA is specifically designed to address the challenges of medical time series: First, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE), which extends the standard RoPE [DBLP:journals/ijon/SuALPBL24] to handle time values using learnable frequency adjustments, allowing MIRA to accurately model the varying frequency intervals of irregularly sampled data. Second, to further improve its ability to adapt to different time features in medical signals, MIRA uses a frequency-specific mixture-of-experts (MoE) layer that learns to route frequency-related information through specialized components. Finally, MIRA integrates a Continuous Dynamics Extrapolation Block, using Neural Ordinary Differential Equations (Neural ODEs) [chen2018neural]. This feature allows the model to predict patient health paths at any future time point, moving beyond the limits of fixed observation grids. MIRA is pretrained on a large curated collection of medical time series, comprised of over 454 billion time points from publicly available datasets. These datasets cover readings from intensive care units (ICUs), operating rooms, pediatric critical care, and long-term sleep and mental health monitoring. All data undergo preprocessing that standardizes time alignment, normalizes sampling frequencies, and maintains clinical realism. MIRA outperforms other time series foundation models with a similar number of activated parameters across various real-world benchmarks, achieving reductions in forecasting errors by an average of 8% and 6% in out-of-distribution and in-distribution scenarios, respectively.
@@ 2 Related Work
Medical Time Series Models. Machine learning has enabled significant progress in healthcare time series analysis, supporting a wide range of tasks, including dynamic forecasting [DBLP:conf/icml/FerteDHGJTLH24, DBLP:conf/ijcai/ChenYW024, DBLP:conf/aaai/Rubin-FalconeLW23, li2025bridge], survival analysis [DBLP:conf/www/Jiang0TY0SCS23, DBLP:conf/ijcai/TanYWY21], clustering and phenotyping [DBLP:conf/aistats/QinSL23, DBLP:conf/aaai/ChenKS22, zhang2023improving], screening and monitoring [DBLP:conf/ijcai/JiaYZDJ24, DBLP:conf/icdm/LiXKFAA022, DBLP:conf/ijcai/YangZC21], early diagnosis [DBLP:conf/aaai/ZhangZL24, DBLP:conf/ijcai/SohnPPC24, DBLP:conf/aaai/HoA23], pharmacology [DBLP:conf/icdm/ZhangLZZZW21], treatment effect estimation [DBLP:conf/kdd/Ye0CA00LWL23, DBLP:conf/aaai/CaoEWSMNL23, li2025bridge], epidemiological and pandemic influenza surveillance [DBLP:journals/ploscb/McIverB14], and hospital resource allocation [DBLP:journals/titb/WangYPEWHGW22, burger2024towards]. Despite this progress, most existing methods are narrowly designed for specific tasks or datasets, limiting their adaptability to diverse clinical scenarios [DBLP:journals/fdata/KaushikCSDNPD20]. Recently, foundation models for medical time series have emerged, showing promising cross-dataset generalization [DBLP:journals/corr/abs-2410-03794, DBLP:conf/icml/ThapaHKMGM024, DBLP:journals/npjdm/WornowXTPSFPFS23]. However, these models focus primarily on classification tasks and are not designed to handle continuous-time forecasting or irregular temporal patterns common in real-world clinical data.
Generalist Time Series Foundation Models. Recent foundation models have shown strong zero-shot forecasting without task-specific tuning [jin2023time, gruver2023large, liu2024unitime, nie2022time, liu2024autotimes]. Chronos [DBLP:journals/corr/abs-2403-07815] first adapted T5 [DBLP:journals/jmlr/RaffelSRLNMZLL20] by discretizing time series intervals as text tokens. More recent models—Moirai [DBLP:conf/icml/WooLKXSS24], Sundial [DBLP:journals/corr/abs-2502-00816], and TimesFM [DBLP:conf/icml/DasKSZ24]—improve support for variable dimensions and frequencies. Time-MoE [DBLP:journals/corr/abs-2409-16040] and Moirai-MoE [DBLP:journals/corr/abs-2410-10469] further improve specialization with sparse experts. However, these models largely require regularly sampled data, limiting their use on irregular clinical data. See Appendix A for more information.
Irregular Time Series Models. Modeling irregular time series has attracted increasing attention, particularly for applications requiring continuous-time reasoning. Early works such as Neural ODEs [chen2018neural, kidger2022neural] and State Space Models (SSMs)[gu2021efficiently, smith2022simplified] have demonstrated the ability to capture continuous dynamics by parameterizing the evolution of latent states over time. Another line of research focuses on adapting deep neural architectures, such as RNNs [DBLP:journals/corr/abs-1907-03907, sun2020review, schirmer2022modeling] and Transformers [DBLP:conf/nips/ChenRWFSL23, ma2022non, xu2019self, zhang2019attain], to irregular settings. While these models have shown success in tasks like interpolation and classification, their effectiveness for long-term forecasting remains underexplored.
@@ 3.1 Irregular Medical Time Series Forecasting
Let each time series instance be represented as paired sequences of timestamps and observations: $\{(t_{i},x_{i})\}_{i=1}^{N}$ where $t_{i}\in\mathbb{R}^{+}$ are strictly increasing ($t_{1}<\cdots<t_{N}$) and $x_{i}\in\mathbb{R}\cup\{\text{NaN}\}$ contains real values or missing entries. Our model is designed to handle two prevalent irregular patterns: (1) Regular-grid missing values. Observations occur at equidistant timestamps $t_{i}=i\Delta t$ but with time-level missing values: some timestamps may have missing measurements. Formally, there exists a subset $\mathcal{M}\subseteq{1,\ldots,N}$ such that the corresponding observations $x_{i}$ are missing (i.e., $x_{i}=\text{NaN}$) for all $i\in\mathcal{M}$, while the timestamps $t_{i}$} remain fully observed. (2) Irregular sampling. Timestamps $t_{i}$ follow non-uniform intervals with $\Delta t_{i}=t_{i+1}-t_{i}$ varying for different $i$. All observations at existing timestamps are present ($x_{i}\in\mathbb{R}$), but the temporal spacing is irregular. Adhering to the channel-independent setting [nie2022time], we formulate the problem in a uni-variate time series manner [DBLP:conf/icml/WooLKXSS24]. Given history $\{(x_{i},t_{i})\}_{i=1}^{L}$, the forecasting task requires predicting future values at known target timestamps $\{t_{L+1},...,t_{L+H}\}$:
@@ 3.2.1 Continuous-Time Rotary Positional Encoding
Standard RoPE [DBLP:journals/ijon/SuALPBL24] assumes discrete and uniformly spaced token indices, which limits its applicability to clinical time series with real-valued and irregular timestamps. To address this, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE) that generalizes RoPE to operate directly on continuous-time inputs. Let $t\geq 0$ denote the timestamp of a token in an irregularly sampled sequence, and let $d$ be the model dimensionality, assumed to be even. Specifically, we discretize continuous timestamps $t\geq 0$ into rotation angles without assuming fixed intervals, enabling the model to handle irregular sampling (more detail can be found at Appendix B). The angular frequencies $\{\omega_{i}\}_{i=0}^{d/2-1}$ are defined as: $\omega_{i}=10000^{-2i/d}$. The resulting time-dependent rotation angle is given by:
@@
This implies that the positional interaction between two tokens is solely a function of their relative timestamp offset (nm)(n-m). In CT-RoPE, this property generalizes to real-valued timestamps, enabling the attention mechanism to capture continuous-time relational structure while retaining the efficiency and structure of standard dot-product attention. In line with DBLP:journals/jmlr/ChowdheryNDBMRBCSGSSTMRBTSPRDHPBAI23, we remove biases except in QKV projections to enhance extrapolation. More detail can be found in Appendix C.
@@ 3.2.3 Continuous Dynamics Extrapolation Block
Auto-regressive transformer architectures, which generate predictions in a stepwise manner under causal masking, cannot incorporate the timestamp of the target token during inference, as it is not available until after generation. To address this limitation and enable extrapolation to arbitrary timestamps, we introduce a Neural ODE-based [chen2018neural] extrapolation module that evolves the latent state from the current token’s timestamp to the target prediction token timestamp, allowing time-aware forecasting at unseen or irregular time points.
@@ 3.3 Model Training
To support generalizable and clinically relevant time series modeling, we curate a large pretraining corpus spanning over 454 billion time points from various real-world healthcare settings. The dataset collection includes signals from ICUs a nd operating rooms, pediatric critical care, long-term sleep and mental health monitoring, and population-level epidemiological surveillance. All data are drawn from publicly available clinical datasets, including MIMIC-III [johnson2016mimic], MIMIC-IV [johnson2020mimic], PTB-XL [wagner2020ptb], Sleep-EDF [kemp2000analysis], and the WAVES Pediatric Waveform Database [miller2023waves]. To enable the model to acquire general knowledge, we resumed training from the Time-MoE checkpoint. A full summary of included datasets and the pre-processing applied is provided in Appendix E.
@@
We adopt a similar model configuration strategy following Time-MoE [DBLP:journals/corr/abs-2409-16040] by providing three model variants of increasing scale: $MIRA_{small}$, with approximately 73 million parameters; $MIRA_{base}$, with approximately 114 million parameters; $MIRA_{large}$, with approximately 455 million parameters;All models are trained on max to eight NVIDIA 80G H/A100 GPUs, using a micro batch size of 128 , and a maximum sequence length of 512. For model configurations and training details, refer to Appendix F.
@@ 4.1 Experiment Setup
Downstream Clinical Datasets. We evaluate MIRA on a diverse suite of real-world clinical and public health datasets, spanning multi-modal physiological signal analysis, critical care monitoring, ambulatory bio-signal tracking, epidemiological surveillance, and healthcare resource utilization. To systematically study temporal robustness, we group datasets into two categories: (1) inherently irregular datasets, which exhibit irregular sampling and natural missing values due to clinical or observational workflows. These are CinC 2012 [silva2012predicting] and CinC 2019 [reyna2020early]. Furthermore, we use (2) originally regular datasets, i.e.,MIT-BIH [moody2001impact], Johns Hopkins COVID-19 Dataset [dong2022johns] ,CDC Influenza Hospitalizations Admissions (CDC-IHA)  [^1] https://www.cdc.gov/flu-forecasting/data-vis/current-week.html, Heart Rate [tan2020monash] and illness [DBLP:conf/iclr/WuHLZ0L23], for which we simulate irregularity by randomly masking 30% of time points. The full list of datasets and statistics is provided in Appendix G.
Baselines. We compare MIRA against 13 state-of-the-art forecasting models, which we categorize into three groups: (i) zero-shot foundation models, including Time-MoE [DBLP:journals/corr/abs-2409-16040]  [^2] We excluded Time-MoE Ultra and Sundial from baselines as they are not open-sourced., Moirai [DBLP:conf/icml/WooLKXSS24], Moirai-MoE [DBLP:journals/corr/abs-2410-10469], Moment [DBLP:conf/icml/GoswamiSCCLD24], TimeGPT [DBLP:journals/corr/abs-2310-03589], Timer [DBLP:conf/icml/LiuZLH0L24], Lag-Llama [rasul2023lag], TimesFM [DBLP:conf/icml/DasKSZ24], and Chronos [DBLP:journals/corr/abs-2403-07815], which require inputs to be interpolated for evaluation; and (ii) full-shot forecasting models, including ContiFormer [DBLP:conf/nips/ChenRWFSL23], T-PatchGNN [DBLP:conf/icml/ZhangYL0024], Neural-CDE [DBLP:conf/nips/RubanovaCD19], and ODE-RNN [DBLP:journals/corr/abs-1907-03907], which are specialized for irregularly sampled time series and require task-specific training. (iii) Continue pre-trained zero-shot foundation models on medical corpora, including Time-MoE [DBLP:journals/corr/abs-2409-16040], Moirai [DBLP:conf/icml/WooLKXSS24] and Chronos [DBLP:journals/corr/abs-2403-07815] which performance best on zero-shot evaluation. Implementation details for all baselines are provided in Appendix H.